OpenStudy (anonymous):

Find an equation of the tangent line to the graph of the function square root(x-2) at the point (18,4)

OpenStudy (anonymous):

Start by finding the derivative.

OpenStudy (anonymous):

Do you know how to do that?

OpenStudy (anonymous):

yes but i must be doing my work wrong

OpenStudy (anonymous):

What are you getting?

OpenStudy (anonymous):

y=4/8x+16

OpenStudy (anonymous):

... You are finding the derivative of \[\sqrt{x-2}\] correct?

OpenStudy (anonymous):

Yes. I got 4/8 for the derivative and than I plugged it into this formula y-y1=m(x-x1)

OpenStudy (anonymous):

Did you get \[\frac{ 1 }{ 2\sqrt{x-2}}\] for the derivative?

OpenStudy (anonymous):

Before you plugged in 18

OpenStudy (anonymous):

yes I did

OpenStudy (anonymous):

Then you should've gotten 1/8, not 4/8?

OpenStudy (anonymous):

oh okay i see what happened. Ok so that is the derivative which also equals the slope?

OpenStudy (anonymous):

Yes. Are you getting the right answer now?

OpenStudy (anonymous):

no. y-4=(1/8)(x-18) y-4=1/8x+9=4 y=1/8x+25/4

OpenStudy (anonymous):

Where did the 9 come from in your second line? 18/8 =/= 9

OpenStudy (anonymous):

Was that supposed to be a slash?

OpenStudy (anonymous):

yes sorry

OpenStudy (anonymous):

Alright, what does your book say the answer is supposed to be?

OpenStudy (anonymous):

Also, you seem to have written 9/4 as positive instead of negative.

OpenStudy (anonymous):

it is an online quiz the answers are: y=x/8+9/4 y=x/8+7/4 y=x/8+9/2 y=x/4+9/2 y=x/4+7/2

OpenStudy (anonymous):

Yes, the answer is the second one. If you look at your work, you'll see that a negative's become a positive, which messed up your answer.

OpenStudy (anonymous):

I see now thank you so much! :)

OpenStudy (anonymous):

Np. :) Since this is apparently your first question, I'll ask you straight-out for a medal, if you understood my explanation.